# Materials Map Integrating Experimental and Computational Data through Graph-Based Machine Learning for Enhanced Materials Discovery

### v2

@@ Abstract
Materials informatics (MI), which emerges from the integration of materials science and data science, is expected to greatly streamline the material discovery and development. The data used for MI are obtained from both computational and experimental studies, while their integration remains challenging. In our previous study, we reported the integration of these datasets by applying a machine learning model that captures trends hidden in the experimental datasets to compositional data stored in the computational database. In this study, we use the obtained data to construct materials maps, which visualize the relation in the structural features of materials, aiming to support study by the experimental researchers. The map is constructed using the MatDeepLearn (MDL) framework, which implements the graph-based representation of material structures, deep learning, and dimensional reduction for the map construction. We evaluate the obtained materials maps through statistical analysis and found that the MDL using message passing neural network (MPNN) enables efficient extraction of features that reflect the structural complexity of materials. Moreover, we found that this advantage does not necessarily translate into improved accuracy in predicting material properties. We attribute this unexpected outcome to the high learning performance inherent in MPNN, which can contribute to the structuring of data points within the materials map.
I Introduction
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of artificial intelligence (AI) and machine learning (ML), researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project (MP) [6] and AFLOW [7] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [8]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications. While computational databases provide extensive datasets derived from first-principles calculations, experimental data remains sparse, inconsistent, and often lacks the structural information necessary for advanced modeling. This disparity between computational and experimental data presents a significant obstacle to the full realization of materials informatics potential.
@@ I Introduction
In general, materials are described by their atomic compositions; however, their properties are fundamentally determined by not only atomic composition but also structural arrangement. To address the limitations of simple chemical formulas in capturing structural nuances, researchers have developed graph-based approaches. The Crystal Graph Convolutional Neural Network (CGCNN) [20] models materials as graphs, where nodes correspond to atoms and edges represent interactions. This method encodes structural information into high-dimensional feature vectors, which, when combined with deep learning techniques, enable the development of robust property prediction models [8].
@@
The implementation of graph-based representations and property prediction models requires advanced computational tools. MatDeepLearn (MDL) [8], developed by Fung et al., provides an efficient Python-based environment for developing material property models. MDL supports various models, including CGCNN [20], Message Passing Neural Networks (MPNN) [21], MatErials Graph Network (MEGNet) [22], SchNet [23], and Graph Convolutional Networks (GCN). The open-source nature and extensibility of MDL make it valuable for researchers implementing graph-based materials property predictions by deep-learning based architectures [8].
@@
The application of graph-based representation materials properties relies heavily on structural data availability [20]. While first-principles calculations inherently include structural information, experimental datasets often lack such details. This limitation poses challenges for applying graph-based methods to experimental data, causing most studies that employ graph-based representations of material properties to rely primarily on computational datasets [8].
@@
StarryData2 (SD2) is an innovative database project in the field of materials science [24]. Developed by Katsura , this database aims to systematically collect, organize, and publish experimental data on materials from previously published papers. SD2 has extracted information from over 7,000 papers, including thermoelectric property data for more than 40,000 samples. The database covers a wide range of material fields, including thermoelectric materials, magnetic materials, low thermal conductivity materials, aggregates, high thermal conductivity materials, and piezoelectric materials. Designed for open access, SD2 allows researchers and engineers to freely access and utilize the data.
@@
In our previous study [25], we proposed a novel framework to integrate experimental and computational datasets to create a comprehensive data for materials informatics. As briefly described in SI, our approach consisted of the following steps: preprocessing experimental data, training a machine learning model, and applying the trained model to predict experimental data for the compositions registered in the computational data. As a result, we obtained a dataset of the predicted experimental values of material properties () with structural information of the corresponding materials.
@@
We demonstrate the effectiveness of the MPNN architecture [21] in capturing structural complexity of materials.
@@ II Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based resentation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [8]. We use the dataset extracted by our previous study, described in Ref. [25] and briefly summarized in SI. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [8], which provides a flexible environment for material property prediction. As schematically described in Figure 1, the model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models, including deep learning approaches. MDL supports the construction of material maps using t-SNE with the data from the first dense layer, which is placed after GC layer as schematically described in Figure 2. This configuration allows the model to capture both local and long-range structural correlations learned by the GC layer.
In this study, we mainly discuss materials maps generated using MPNN [21], which produces well-structured materials map efficiently. In MDL, MPNN is characterized by the Graph Convolutional (GC) layer configured by the neural network (NN) layer and the gated recurrent unit (GRU) layer. The NN layer enhances the model’s representational capacity, while the GRU layer improves learning efficiency through memory mechanisms. The number of GC layer repetitions was set to 4, which is the default hyperparameter [8] of MDL, unless otherwise specified.
III Results
III.1 Material Map Generated Using MPNN Architecture
@@
Let us first demonstrate a material map generated using the MPNN [21] architecture in Figure 2. The default hyperparameters of MDL [8] were used in the model training. The x and y axes represent the first and second t-SNE components, while the color of each point represents the predicted experimental value. In the figure, we find a clear trend where lower values concentrate in the lower region, while higher values appear in the upper region. The smooth color gradients reflect meaningful relationships between predicted values and structural features. We also find two distinct branches, labeled BR1 and BR2 in Figure 2, extended laterally in the upper portion, featuring significant clusters of high values at their endpoints. The fine structures visible in the map suggest that the model effectively captures structural features of materials.
III.2 Contribution of GC Layer in MPNN Architecture for Map Construction
@@
In MDL, the graph-based architecture is implemented in the GC layer, characterized by GC type, GC dimensions, and the repetition number of GC blocks (N). To investigate the role of the GC layer, we compare the maps generated by MDL with N = 1 and 10 in Figure 3. We find that the increase in N leads to tighter clustering of data points. To quantify the structures in the map, we evaluate the distributions of the nearest neighbor distances (NND) of each data point by Kernel Density Estimation (KDE) [26]. KDE is a non-parametric method for estimating smooth probability density functions from data points, and thus offers a more continuous and less biased distribution representation than histograms. Figure 3(c) summarizes the KDE spectra obtained with N = 1, 4 (default value for MDL), and 10. With increasing N, the peak intensity increases and the peak position shifts to the lower NND region, convincing the enhanced feature learning by the GC layer.
@@
The framework of DeeperGATGNN allows machine learning modeling with large N [15]. With increasing N, prediction accuracy and structural feature learning are enhanced, while memory usage for computation increases dramatically especially when large datasets are used for the analysis [15]. Therefore, the use of MDL with MPNN architecture may be limited to cases where the dataset size is relatively small or when computing resources with sufficient memory size are available.
@@
In MDL, the MPNN architecture is characterized by NN and GRU blocks in the GC layer. To investigate the contributions of each block, we selectively controlled the training of each block, turning them ON or OFF, by modifying the MDL code. N was again set to 10 to enhance the contribution of each block. Figures 4(a)-4(d) illustrate the effects of disabling learning in either the NN or GRU blocks. When learning is disabled in either block, the points spread more diffusely; when both are disabled, the points distribute uniformly across the map. These trends are again clearly captured by the results of KDE, shown in Figure 4(e), as the increase in the peak intensity and the peak shift toward the lower NND region by turning ON the training by either or both of the NN and GRU blocks. Overall, these results demonstrate that learning by the GC layer, composed of the NN and GRU blocks, plays a crucial role the structural organization in the materials map.
III.3 Model Dependency of Material Maps
@@
So far, we have discussed material maps generated by MDL using the MPNN architecture. MDL also supports other architectures, which are highly sophisticated and specialized for analyzing material properties, such as CGCNN [20], MEGNet [22], GCN, and SchNet [21]. In Figures 5(a)-5(d), we compare the material maps generated by each of these models and find that the maps generated by the other models lack the clear structures observed in the map generated by MPNN. These trends are again evaluated by the distribution analysis of NND of each data point by KDE, as summarized in Figure 5(e). We find that the KDE spectrum for the MPNN-generated map displays a higher density in the low NND region compared to the spectra for other models. These trends are emphasized by increasing the repetition number the GC layer, leading us to convince the superiority of MPNN as an effective tool providing more structurally organized and continuous mappings compared to other graph-based models.
@@
Similar discussions are found in Ref. [8], which is the first report of MDL. This study also compares material maps generated with different models; however, the differences in the obtained maps were not as significant compared to the present case. We attribute the difference in these two studies to the data screening applied in our study. Both studies use the same database, Materials Project, while, in this study, we applied a filter based on material stability, band gap, and atomic elements. This screening results in the selection of similar materials, which facilitated feature extraction by machine learning models, leading to material maps with clearer structures.
IV Discussion
IV.1 Relationship between material property prediction and material map construction
@@
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [25].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R2), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset, summarized in Table 1, leads that, while the MPNN produces the most structurally interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result leads us to conclude that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@
It may be worth noting that MPNN is a graph-based architecture developed relatively early in the field [21] and is characterized by high learnability and versatility. Due to the black-box nature of deep learning-based modeling, it is unclear why the MPNN-generated materials map can capture materials properties so efficiently. Further investigation into the reasons behind these discrepancies could yield valuable insights into how structural interpretability and predictive power might be balanced in future model designs.
IV.2 Improvement of the interpretability of material maps
So far, we mainly discussed the materials maps colored by the predicted values, while the computational dataset obtained from the Materials Project encompasses diverse materials data. To improve the interpretability of the obtained maps, we made materials maps colored by the other properties registered in the Materials Projects. Figure 6 shows the maps colored by (a) predicted value, (b) energy per atom, (c) number of sites, (d) volume, and (e) number of elements. The following analysis using these rich computational datasets with materials maps provides detailed insights that contribute to a more comprehensive interpretation of the results.
In Figure 6(a), colored by predicted , we find the lower values in the bottom left side while high values in the top right and top left sides. Similar trends are found in Figure 6(b), colored by the energy per atom, representing the close relationships between the corresponding parameters. This trend is supported by the high correlation coefficient of 0.66, obtained through the correlation analysis described in SI.
@@ IV.2 Improvement of the interpretability of material maps
To investigate the distribution of materials in the map, we applied -means clustering analysis to the map with =10 clusters as shown in Figure 6(f). The label for each cluster was assigned by sorting based on the average of predicted values, allowing us to identify regions of high-performance materials. To understand the materials included in each cluster, we counted the number of elements in the compositions of the materials included in each cluster. The elements with less than 10 of the compositional ratio was excluded to account for the elements that make a major contribution to the material properties. The results for each cluster are plotted using the Periodic Trend Plotter (https://github.com/Andrew-S-Rosen/periodic_trends) in Figure 7. In each cluster, we find distinct characteristic distributions of S, Se, and Te atoms. Clusters containing a large number of materials with low zT values (1–7) show a high frequency of S and Se occurrences. On the other hand, clusters containing a large number of materials with high zT values (8–10) exhibit a higher frequency of Te occurrences. In order to find the characteristics of the two branches, labeled as BR1 and BR2 in Figure 2, we compare the elemental distribution for clusters 9 and 10, which are placed at the edge of each branch. We find high similarities between these two plots, such as the high frequencies of Te , Sb, Ti, and Ag atoms. The detailed discussion in terms of the atomic distributions supports our improved understanding of the obtained material maps.
IV.3 Finding proper materials with material maps

### v3

@@ Abstract
Materials informatics (MI), which emerges from the integration of materials science and data science, is expected to greatly streamline material discovery and development. The data used for MI are obtained from both computational and experimental studies, while their integration remains challenging. In our previous study, we reported the integration of these datasets by applying a machine learning model that captures trends hidden in the experimental datasets to compositional data stored in the computational database. In this study, we use the obtained data to construct materials maps, which visualize the relation in the structural features of materials, aiming to support study by the experimental researchers. The map is constructed using a MatDeepLearn (MDL) framework, which implements the graph-based representation of material structures, deep learning, and dimensional reduction for map construction. We evaluate the obtained materials maps through statistical analysis and found that MDL using message passing neural network (MPNN) architecture enables efficient extraction of features that reflect the structural complexity of materials. Moreover, we found that this advantage does not necessarily translate into improved accuracy in the prediction of material properties. We assume this unexpected outcome to the high learning performance inherent in MPNN, which can contribute to the structuring of data points within the materials map.
keywords:
Materials informatics; Visualization; Material map; Graph theory; Deep learning1 Introduction
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of machine-learning and/or deep-learning techniques, researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods [6, 7, 8, 9, 10, 11, 12]. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project (MP) [13] and AFLOW [14] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [15]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications.
@@
In general, materials are described by their atomic compositions; however, their properties are fundamentally determined by not only atomic composition but also structural arrangement. To address the limitations of simple chemical formulas in capturing structural nuances, researchers have developed graph-based approaches. The Crystal Graph Convolutional Neural Network (CGCNN) models materials as graphs, where nodes correspond to atoms and edges represent interactions [27]. This method encodes structural information into high-dimensional feature vectors, which, when combined with deep learning techniques, enable the development of robust property prediction models [15].
@@
The implementation of graph-based representations and property prediction models requires advanced computational tools. MatDeepLearn (MDL), developed by Fung ., provides an efficient Python-based environment for developing material property models [15]. MDL supports various models for graph-based representation of materials structures, including CGCNN [27], Message Passing Neural Networks (MPNN) [28], MatErials Graph Network (MEGNet) [29], SchNet [30], and Graph Convolutional Networks (GCN). The open-source nature and extensibility of MDL make it valuable for researchers implementing graph-based materials property predictions with deep-learning based architectures [15].
@@
The application of graph-based representation of materials properties relies heavily on structural data availability [27]. While computational databases provide extensive datasets derived from first-principles calculations, experimental data remains sparse, inconsistent, and often lacks the structural information necessary for advanced modeling. This limitation poses challenges for applying graph-based methods to experimental data, causing most studies that employ graph-based representations of material properties to rely primarily on computational datasets [15].
@@
StarryData2 (SD2) is an innovative database project in the field of materials science [31]. Developed by Katsura ., this database aims to systematically collect, organize, and publish experimental data on materials from previously published papers. SD2 has extracted information from over 7,000 papers for more than 40,000 samples. The database covers a wide range of material fields, including thermoelectric materials, magnetic materials, thermal conductivity materials, and piezoelectric materials. Designed for open access, SD2 allows researchers and engineers to freely access and utilize the data.
@@
In our previous study [32], we proposed a novel framework to integrate experimental and computational datasets to create a comprehensive data for materials informatics. As briefly described in Appendix A, our approach consisted of the following steps: preprocessing experimental data, training a machine learning model, and applying the trained model to predict experimental data for the compositions registered in the computational data. As a result, we obtained a dataset of the predicted experimental values of material properties () with structural information of the corresponding materials.
@@
We demonstrate the effectiveness of the MPNN architecture in capturing structural complexity for the material map construction.
@@ 2 Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based representation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [15]. We use the dataset extracted by our previous study, described in Ref. [32] and briefly summarized in Appendix A. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [15], which provides a flexible environment for material property prediction. The model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models. MDL supports the construction of material maps using t-SNE with the data from the first dense layer, which is placed after GC layer as schematically described in Figure 1.
3 Results
3.1 Material Map Generated Using MPNN Architecture
@@
Let us first demonstrate a material map generated using the MPNN [28] architecture in Figure 2. The default hyperparameters of MDL [15] were used in the model training. The x and y axes represent the first and second t-SNE components, while the color of each point represents the predicted experimental value. In the figure, we find a clear trend where lower values concentrate in the lower region, while higher values appear in the upper region. The smooth color gradients reflect meaningful relationships between predicted values and structural features. Moreover, the fine structures visible in the map suggest that the model effectively captures structural features of materials. We also find two distinct branches, labeled BR1 and BR2 in Figure 2, extended laterally in the upper portion, featuring significant clusters of high values at their endpoints.
3.2 Contribution of GC Layer in MPNN Architecture for Map Construction
@@
In MDL, MPNN is defined by the Graph Convolutional (GC) layer configured by a neural network (NN) layer and a gated recurrent unit (GRU) layer. The NN layer enhances the model’s representational capacity, while the GRU layer improves learning efficiency through memory mechanisms. The GC layer is characterized by GC type, GC dimensions, and the repetition number of GC blocks (N). In this study, N was set to 4, which is the default hyperparameter of MDL [15], unless otherwise specified. To investigate the role of the GC layer, we compare maps generated by MDL with N = 1 and 10 in Figure 3. We find that the increase in N leads to tighter clustering of data points. To quantify the structures in the map, we evaluate the distributions of the nearest neighbor distances (NND) of each data point by Kernel Density Estimation (KDE) [33]. KDE is a non-parametric method for estimating smooth probability density functions from data points, and thus offers a more continuous and less biased distribution representation than histograms. Figure 3(c) summarizes the KDE spectra obtained with N = 1, 4 (default value for MDL), and 10. With increasing N, the peak intensity increases and the peak position shifts to the lower NND region, convincing the enhanced feature learning by the GC layer.
@@
With increasing N, prediction accuracy and structural feature learning are enhanced, while the memory usage for computation increases dramatically especially when large datasets are used for the analysis [22]. This is demonstrated by the study using DeeperGATGNN, which allows modeling with large N [22]. Therefore, the use of MDL with MPNN architecture may be limited to cases where the dataset size is relatively small or when computing resources with sufficient memory size are available.
@@
In MDL, the MPNN architecture is characterized by NN and GRU blocks in the GC layer. To investigate the contributions of each block, we selectively controlled the training of each block, turning them ON or OFF, by modifying the MDL code. N was again set to 10 to enhance the contribution of each block. Figures 4(a)-4(d) illustrate the effects of disabling learning in either the NN or GRU blocks. When learning is disabled in either block, the points spread more diffusely; when both are disabled, the points distribute uniformly across the map. These trends are again clearly captured by the results of KDE, shown in Figure 4(e). By turning ON the training by either or both of the NN and GRU blocks, the peak intensity increases and the peak position shifts toward the lower NND region. These results convince us that the learning by the GC layer, composed of the NN and GRU blocks, plays a crucial role in the structural organization in the materials map.
3.3 Model Dependency of Material Maps
@@
So far, we have discussed material maps generated by MDL using the MPNN architecture. MDL also supports other architectures, such as CGCNN [27], MEGNet [29], GCN, and SchNet [28]. In Figures 5(a)-5(d), we compare the material maps generated by each of these models and find that the maps generated by the other models lack the clear structures observed in the map generated by MPNN. These trends are again evaluated by the distribution analysis of NND of each data point by KDE, as summarized in Figure 5(e). We find that the KDE spectrum for the MPNN-generated map displays a higher density in the low NND region compared to the spectra for other models. These trends are emphasized by increasing the repetition number the GC layer, leading us to convince the superiority of MPNN as an effective tool providing more structurally organized and continuous mappings compared to other graph-based models.
@@
Similar discussions are found in Ref. [15], which is the first report of MDL. This study also compares material maps generated with different models; however, the differences in the obtained maps were not as significant compared to the present case. We attribute the difference in these two studies to the data screening applied in our study. Both studies use the same database, Materials Project, while we here applied a filter based on material stability, band gap, and atomic elements. This screening results in the selection of similar materials, which facilitated feature extraction by the machine learning, leading to material maps with clearer structures.
4 Discussion
4.1 Relationship between material property prediction and material map construction
@@
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [32].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R2), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset are summarized in Table 1. While the MPNN produces the most structurally interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result leads us to conclude that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@
MPNN is a graph-based architecture developed relatively early in the field [28] and is characterized by high learnability and versatility. Due to the black-box nature of deep learning-based modeling, it is unclear why the MPNN-generated materials map can capture materials properties so efficiently. Further investigation into the reasons behind these discrepancies could yield valuable insights into how structural interpretability and predictive power might be balanced in future model designs.
4.2 Improvement of the interpretability of material maps
So far, we mainly discussed the materials maps colored by the predicted values, while the computational dataset obtained from the Materials Project encompasses diverse materials data. To improve the interpretability of the obtained maps, we made materials maps colored by the other properties. Figure 6 shows the maps colored by (a) predicted value, (b) energy per atom, (c) number of sites, (d) volume, and (e) number of elements. The following analysis using these rich computational datasets with materials maps provides detailed insights of the obtained maps.
In Figure 6(a), colored by predicted , we find the lower values in the bottom left side while high values in the top right and top left sides. Similar trends are found in Figure 6(b), colored by the energy per atom, representing the close relationships between the corresponding parameters. This trend is supported by the high correlation coefficient of 0.66, obtained through the correlation analysis described in Appendix B.
@@ 4.2 Improvement of the interpretability of material maps
To investigate the distribution of materials in the map, we applied -means clustering analysis to the map with =10 clusters as shown in Figure 6(f). The label for each cluster was assigned by sorting based on the average of predicted values, allowing us to identify regions of high-performance materials. To understand the materials included in each cluster, we counted the frequency of elements in the compositions of materials within each cluster. The elements with less than 10 of the compositional ratio were excluded to focus on elements that make major contributions to the material properties. The results for each cluster are plotted using the Periodic Trend Plotter (https://github.com/Andrew-S-Rosen/periodic_trends) in Figure 7. In each cluster, we find distinct characteristic distributions of S, Se, and Te atoms. Clusters 1–7 containing materials with low values show a high frequency of S and Se occurrences. On the other hand, clusters 8–10 containing materials with high values exhibit a higher frequency of Te occurrences. In order to find the characteristics of the two branches, labeled as BR1 and BR2 in Figure 2, we compare the elemental distribution for clusters 9 and 10, which are placed at the edge of each branch. We find high similarities between these two plots, such as the high frequencies of Te, Sb, Ti, and Ag atoms. Namely, the materials in clusters 9 and 10 have similar compositions, but are separated in the map by their structural complexity.
4.3 Finding proper materials with material maps